OpenStudy (anonymous):

Some please tell me what you get for an answer! Solve by the quadratic formula: 4x^2-3x+1=0

OpenStudy (anonymous):

if u notice it, the discriminant is less than zero, so it has no real solutions

OpenStudy (anonymous):

So what should I put? And what's "discriminant"?

myininaya (myininaya):

\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}, ax^2+bx+c=0\]

OpenStudy (anonymous):

meaning the b^2-4ac part is less than zero

OpenStudy (anonymous):

I know how to do it myininaya, I just got confusing answers

OpenStudy (anonymous):

therefore, you can't take the square root of it

myininaya (myininaya):

\[a \neq 0\]

myininaya (myininaya):

what is a what is b what is c

OpenStudy (anonymous):

So what would I write for an answer? Not possible?

myininaya (myininaya):

what is a what is b what is c

myininaya (myininaya):

i will help you if you answer my questions

OpenStudy (anonymous):

Ok, Ok, a=4 b=-3 c=1

myininaya (myininaya):

\[\frac{-(-3) \pm \sqrt{(-3)^2-4(4)(1)}}{2(4)}\]

myininaya (myininaya):

\[\frac{3 \pm \sqrt{9-16}}{8}=\frac{3 \pm \sqrt{-7}}{8}\]

myininaya (myininaya):

\[=\frac{3 \pm i \sqrt{7}}{8}\]

OpenStudy (anonymous):

What does i mean?

myininaya (myininaya):

imaginary \[\sqrt{-1}=i\]

myininaya (myininaya):

not real

OpenStudy (anonymous):

So is the third to last thing that you put be the answer on my test Wednesday? Or would it be some other term?